OpenStudy (anonymous):

Pls Help Quick! Solve the following system of equations for x. - 3x + 7y = 2 - 2x - 3y = - 4 A. x=2/23 B.x=22/23 C.x=-2/23 D.x=-22/23

OpenStudy (anonymous):

@Legends @RiahB13 @jagr2713 @satellite73

OpenStudy (anonymous):

@Holly00d1248

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

One Moment

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Have you solved for y?

OpenStudy (anonymous):

no.was i suppose to?

OpenStudy (anonymous):

im not very good at this

OpenStudy (anonymous):

Solving for y will tell you the value for x

OpenStudy (anonymous):

You're fine, we have all struggled with math, we'll work through this problem. :-)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Do you know how to solve for y?

OpenStudy (anonymous):

:l nope

OpenStudy (anonymous):

my school never went over this

OpenStudy (anonymous):

To solve for y we can use elimination method to isolate the y value and eliminate the x value.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

What must be true about x if 3+x=0

OpenStudy (anonymous):

:| i suck at this

OpenStudy (anonymous):

You're fine, I am here to help :) (-3x) (-2x) In order to make them equal to each other what do you think you should do?

OpenStudy (anonymous):

get rid of the x?

OpenStudy (anonymous):

You need to make it so you have a positive number plus a negative number so the cancel out. How would I go about making -3x positive?

OpenStudy (anonymous):

*they

OpenStudy (anonymous):

add 6x to it? im horrible at math XD

OpenStudy (anonymous):

I wish I could help but Im terrible too XD Haha its NOT my thing :'

OpenStudy (anonymous):

You wont use addition or subtraction yet. Hint:You do need to set it equal to 6 :)

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

What another way of representing 3+3?

OpenStudy (anonymous):

umm...

OpenStudy (anonymous):

Have you learned multiplication?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

What does 3+3= ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

3*2=?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

What are the values of x in your problem?

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):

i...dont know

OpenStudy (anonymous):

my school didnt go over this so i dont know alot about it...sorry

OpenStudy (anonymous):

-3 and -2, and as we just covered in order to set them equal to each other we will need to multiply the entire top equation by -3 and the entire bottom equation by 2

OpenStudy (anonymous):

-2 or just 2

OpenStudy (anonymous):

If you multiply the top by a negative you get a positive 6 so you need to keep the negative on the bottom or else you will get positive 6 and 6+6=12 not 0

OpenStudy (anonymous):

Still with me?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

What do you have now?

OpenStudy (anonymous):

sorry my computer is lagging

OpenStudy (anonymous):

o.o i dont know sorry

OpenStudy (anonymous):

@Tylyngaming

OpenStudy (anonymous):

i got it,thanks! ill close the question @Prometheus777